# Impurity immersed in a two-component few-fermion mixture in a one-dimensional harmonic trap

### v1

@@ Abstract
We investigate a one-dimensional three-component few-fermion mixture confined in a parabolic external trap, where one component contains a single particle acting as an impurity. Focusing on the many-body ground state, we analyze how the interactions between the impurity and the other components influence the system’s structure. For fixed interaction strengths within the mixture, we identify a critical interaction strength with the impurity for which the system undergoes a structural transition characterized by substantial change in its spatial features. We explore this transition from the point of view of correlations and ground-state susceptibility. We remarkably find that this transition exhibits unique universality features not previously observed in other systems, highlighting novel many-body properties existing in multi-component fermionic mixtures.
@@ I Introduction
The experimental study of ultracold gases has opened up many new possibilities for discovering previously not well-known quantum phenomena. This is due to the possibilities offered by the rich internal structure of the atoms. In particular, the existence of metastable atomic states with high total spin allows one to study exotic mixtures of quantum particles obeying various quantum statistics and experiencing highly non-trivial couplings between spin and spatial degrees of freedom Cazalilla and Rey (2014); Pagano et al. (2014); Sonderhouse et al. (2020). Their properties in reduced dimensionality have become the subject of intensive research at many levels Liu et al. (2008); Jiang et al. (2009); Guan et al. (2010); Lee et al. (2011); Kuhn and Foerster (2012); Guan et al. (2012); Jiang et al. (2016); Laird et al. (2017); Dobrzyniecki and Sowiński (2020); Silva-Valencia and Mendoza-Arenas (2025) but their collective behavior induced by interactions in the few-body regime still require systematic investigation.
In this work, we analyze a mesoscopic fermionic mixture consisting of three distinguishable components. Our aim is to test whether any collective effects may appear in the case of a larger number of components, which are not present in binary mixtures. In order to make these studies well-defined, reasonably comprehensive, and also going beyond extensively studied balanced mixtures, we consider the smallest possible extension of two-component fermionic mixture by introducing additional interaction with a single third-component particle. We aim to investigate if and how interactions with this impurity change the internal properties of the mixture. Moreover, we inspect non-classical correlations this may induce. In this way we also link our work to the well-studied problem of a single impurity interacting with several spin-polarized fermions Wenz et al. (2013); Włodzyński (2022).
Our work is organized as follows. In Sec. II we introduce the theoretical model of the system studied and we describe its Hamiltonian in the convenient formalism. Then in Sec. III we briefly explain the exact diagonalization method in the energy cut-off framework. In Sec. IV we analyze single- and two-particle density distributions and we identify a specific transition in the many-body ground state of the system induced by interactions with impurity. In Sec. V we study this transition in detail and we show that it has intriguing universal properties that do not depend essentially neither on the number of particles in the system nor the internal interactions in the mixture. Finally, we conclude in Sec. VI.
@@ II The system
In our work, we focus on the simplest possible extension of the two-component mixture of fermions to scenarios where interactions with additional third-component particle (impurity) are present. We assume that all the particles have the same mass $m$, they are confined in the same one-dimensional parabolic trap of frequency $\Omega$, and they mutually interact via zero-range contact interactions in $s$-wave channel. This implies of course that, due to the Pauli exclusion principle, only interactions between particles from different components do not vanish. Since in our considerations the third component is essentially distinguished, we write the Hamiltonian of the system conveniently as
^=^$\hat{\cal H}=\hat{\cal H}_{AB}+\hat{\cal H}_{C}+\hat{\cal H}_{I}.$ (1)
In this sum the Hamiltonian ^$\hat{\cal H}_{AB}$ describes the two-component mixture of fermions and in the second quantization reads
^$\hat{\cal H}_{AB}$ $=\int\!\mathrm{d}x\sum_{\sigma}\,\hat{\Psi}_{\sigma}^{\dagger}(x)\left[-\frac{\hbar^{2}}{2m}\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}+\frac{m\Omega^{2}}{2}x^{2}\right]\hat{\Psi}_{\sigma}(x)$
@@
Here $\sigma\in\{A,B\}$ enumerates mixture’s components while $\hat{\Psi}_{\sigma}(x)$ is a field operator annihilating $\sigma$-component particle at position $x$. These operators obey fermionic anti-comutation relations $\{\Psi_{\sigma}(x),\Psi^{\dagger}_{\sigma^{\prime}}(x^{\prime})\}=\delta_{\sigma\sigma^{\prime}}\delta(x-x^{\prime})$ and $\{\Psi_{\sigma}(x),\Psi_{\sigma^{\prime}}(x^{\prime})\}=0$. Dimensionless parameter $g_{0}$ quantifies an effective interaction strength between components.
Contrary, the Hamiltonian ^$\hat{\cal H}_{C}$ describes a third-component particle confined in the same trap. In the first quantization, it is represented by the differential operator of the form
^$\hat{\cal H}_{C}=-\frac{\hbar^{2}}{2m}\frac{\mathrm{d}^{2}}{\mathrm{d}z^{2}}+\frac{m\Omega^{2}}{2}z^{2},$ (2b)
@@
Finally, the Hamiltonian ^$\hat{\cal H}_{I}$ is responsible for interactions between the two-component mixture and the third particle. It can be written as
^$\hat{\cal H}_{I}=g\sqrt{\frac{\hbar^{3}\Omega}{m}}\int\!\mathrm{d}x\sum_{\sigma}\,\hat{\Psi}^{\dagger}_{\sigma}(x)\delta(x-z)\hat{\Psi}_{\sigma}(x),$ (2c)
@@ III The Method
where $|\phi_{n}\rangle$ is a state of the third-component particle occupying $n$-th single-particle orbital, while two integer vectors $\vec{\alpha}=(\alpha_{1},\ldots,\alpha_{N})$ and $\vec{\beta}=(\beta_{1},\ldots,\beta_{N})$ encode single-particle orbitals occupied by individual fermions from components $A$ and $B$, respectively. Indistinguishability and fermionic nature of particles is guaranteed by imposing additional constrains of the form $\alpha_{1}>\alpha_{2}>\ldots>\alpha_{N}$ and $\beta_{1}>\beta_{2}>\ldots>\beta_{N}$. States (5) are obviously eigenstates of the non-interacting Hamiltonian ($g_{0}=g=0$) and have energies $E_{\vec{\alpha},\vec{\beta},n}=\hbar\Omega\left[(2N+1)/2+n+\sum_{i}(\alpha_{i}+\beta_{i})\right]$. Non-interacting ground state has energy $E_{0}=N^{2}+1/2$.
@@
To diagonalize numerically the Hamiltonian one needs to cut the basis. In our approach, we do not perform simple cuts on single-particle excitations but rather we limit the many-body basis to states having total (non-interacting) energy $E_{\vec{\alpha}\vec{\beta}n}$ not larger than some fixed energy limit $E_{C}$. In this way, we do not exclude from the description low-energy states having only a few particles highly excited, while we neglect high-energy states with many low-excited particles. It was shown in previous studies that such a method gives a much better approximation of interacting many-body states with the same usage of numerical resources Haugset and Haugerud (1998); Płodzień et al. (2018); Chrostowski and Sowiński (2019); Rojo-Francàs et al. (2022). Following this procedure, for a given number of particles $N$ and interaction strengths $g_{0}$ and $g$, we represent the Hamiltonian (1) in the truncated many-body basis and diagonalize with the implicitly-restarted Arnoldi-Lanchos algorithm Lehoucq et al. (1998). As a result we obtain the lowest eigenstates and their energies. Particularly, the many-body ground state $|\mathtt{G}\rangle$ having energy ${\cal E}_{\mathtt{G}}$ is represented as a superposition of basis states
$|\mathtt{G}\rangle=\sum_{\vec{\alpha},\vec{\beta},n}G_{\vec{\alpha}\vec{\beta}n}|{\vec{\alpha},\vec{\beta},n}\rangle.$ (6)
Let us mention here that in the case of harmonic confinement further optimization techniques of the many-body basis are possible Volosniev, A.G. et al. (2015); Bellotti, Filipe F. et al. (2017); Kościk (2018, 2020). They are vital when large interaction strengths are considered. In the cases of relatively weak interactions studied in our work, this kind of optimization has limited importance. Therefore, we do not introduce them to our scheme.
@@ IV Impact of the impurity
Figure 1: Impact of interactions with the third-component particle on the two-component mixture containing four particles ($N=2$). Different columns correspond to different values of interaction strength with the impurity $g$. In all plots, inter-component interaction strength is fixed, $g_{0}=1$. (a) Single-particle density profiles of the mixture’s components $n_{\sigma}(x)$ (thick blue) and the third-component particle $m(x)$ (thin red). For a better comparison, the non-interacting distribution of the mixture’s components is additionally displayed with a light blue shadow. (b) Two-particle density profile $\mu_{\sigma}(x,z)$. (c) Distribution of correlations ${\cal G}_{\sigma}(x,z)$ between impurity and particle from the component $\sigma$. Note, that in the left plot, we consider the interaction strength $g$ slightly positive ($g=0.01$) since for vanishing interactions the distribution vanishes. In this case, the distribution intensity is also scaled by a factor $10^{2}$ to make it visible. In all plots, positions are expressed in units of $a_{0}$. The upper and lower values on the color bar refer to plots (b) and (c), respectively.
Figure 2: Distribution of internal two-particle correlations ${\cal K}(x,y)$ for the system of five particles ($N=2$) for different internal interaction strengths $g_{0}$ and interactions with the impurity $g$. For any fixed interaction $g_{0}$, along with increasing repulsion $g$, the transition between two different regimes is visible. In all plots, positions are expressed in units of $a_{0}$.
@@
where ^$\hat{\mathbb{P}}_{x}=|x\rangle\langle x|$ projects to an eigenstate $|x\rangle$ of position operator. Of course, due to the exchange symmetry $A\leftrightarrow B$ explained before, both density distributions $n_{A}(x)$ and $n_{B}(x)$ are exactly the same. In Fig. 1(a), we present these distributions (thick blue and thin red lines) for the system of five particles ($N=2$) for different interaction strengths $g$, keeping fixed internal interaction $g_{0}=1$. For better comparison, we additionally display with the shadowed area the non-interacting ($g_{0}=g=0$) distribution $n_{\sigma}(x)$. It is clear that along with increasing $g$, the distribution of the impurity $m(x)$ significantly changes its character and for sufficiently large repulsions it is split and pushed out from the center where the mixture is concentrated. In contrast, distributions of mixture components $n_{\sigma}(x)$ are only slightly affected. This behavior of density distributions can be investigated further by considering the two-particle density distribution
@@
which encodes the joint probability of finding the impurity at position $z$ and one of the particles from the component $\sigma$ at position $x$. The distribution for the same parameters of the system is shown in Fig. 1(b). In addition to the spatial separation already deduced from single-particle distributions, the two-particle density clearly indicates that some correlations between mixture and impurity are developed. Most significantly they are visible along the diagonal direction $z=x$, i.e., the probability of finding both particles at the same position quickly vanishes with increasing interaction $g$. Of course, this effect cannot be captured by single-particle distributions and is not present in the direct product $n_{A}(x)m(z)$. To make it evident we additionally display in Fig. 1(c) the distribution of correlations defined as
@@
From the perspective of the internal properties of the mixture, the description is slightly simpler. As noticed already, interactions with the impurity do not influence significantly the single-particle distribution of the mixture. It turns out that this is also the case when one considers the inter-component two-particle distribution
@@
The impact of the impurity is however noticeable when the distribution of two-particle correlations is analyzed. It is defined analogously to correlations (9) as
@@
In Fig. 2 we show this distribution for different values of interactions $g_{0}$ and $g$ for the same system of five particles ($N=2$). The magnitude of correlations is very small when compared to the magnitude of correlations in ${\cal G}_{\sigma}(x,z)$. However, in the region of parameters where the impurity starts to leak outside the mixture, its structural transition is clearly visible.
@@
From phenomenological argumentation, some kind of separation of the components is suspected. When interactions are strong enough the single-particle excitations become energetically favourable over interaction energy cost. Then, to reduce the overlap of the densities the system excites particles. It should be noted, however, that the separation observed here is substantially different from separations observed in systems of single-component fermions interacting with the impurity Lindgren et al. (2014). It was shown that in those systems, the impurity remains in the middle of the trap while the fermionic component is (only slightly) split and pushed out from the center. In our case, although repulsive interactions in the fermionic mixture are present, the distribution of the impurity is split and driven to the edges. As noted in the literature, in the case of the single-component Fermi sea, the existence of this kind of separation requires a noticeable mass difference between impurity and other particles Pȩcak et al. (2016); Włodzyński (2022).
@@
Figure 3: Properties of the fidelity susceptibility for different numbers of particles in the system. (a) The fidelity susceptibility $\chi(g_{0};g)$ as a function of interaction $g$ for different internal interactions $g_{0}$. (b) Position of the transition $G_{N}(g_{0})$ as a function of internal interaction $g_{0}$. (c) Susceptibilities from the left plot after rescaling in magnitude by $\chi_{0}(g_{0})$ and centering around the transition interaction $G_{N}(g_{0})$. All plots collapse to the same universal curve signaling the specific universality of the transition. (d) Numerical collapse to the universal curve $\Phi$ (dashed black) of all the fidelity susceptibility curves (different points and colors) after applying appropriate scaling (15). Note that deviations from the universality are diminished when the number of particles in the mixture increases.

### v2

@@ Abstract
We investigate a one-dimensional three-component few-fermion mixture confined in a parabolic external trap, where one component contains a single particle acting as an impurity. Focusing on the many-body ground state, we analyze how the interactions between the impurity and the other components influence the system’s structure. For fixed interaction strengths within the mixture, we identify a critical interaction strength with the impurity for which the system undergoes a structural transition characterized by a substantial change in its spatial features. We explore this transition from the point of view of correlations and ground-state susceptibility. We remarkably find that this transition exhibits unique universality features not previously observed in other systems, highlighting novel many-body properties existing in multi-component fermionic mixtures.
@@ I Introduction
The experimental study of ultracold gases has opened up many new possibilities for discovering previously not well-known quantum phenomena. This is due to the possibilities offered by the rich internal structure of the atoms. In particular, the existence of metastable atomic states with high total spin allows one to study exotic mixtures of quantum particles obeying various quantum statistics and experiencing highly non-trivial couplings between spin and spatial degrees of freedom Cazalilla and Rey (2014); Pagano et al. (2014); Sonderhouse et al. (2020). Their properties in reduced dimensionality have become the subject of intensive research at many levels Liu et al. (2008); Jiang et al. (2009); Guan et al. (2010); Lee et al. (2011); Kuhn and Foerster (2012); Guan et al. (2012); Jiang et al. (2016); Laird et al. (2017); Dobrzyniecki and Sowiński (2020); Anh-Tai et al. (2024); Silva-Valencia and Mendoza-Arenas (2025); Tran et al. (2025), but their collective behavior induced by interactions in the few-body regime still requires systematic investigation. Moreover, due to the first spectacular experiments on multi-component few-body systems Ottenstein et al. (2008); Schumacher et al. (2023), a deeper theoretical understanding of systems hosting more than two components became very necessary. In our work, we take the first step in this direction.
In this work, we analyze a mesoscopic fermionic mixture consisting of three distinguishable components. Our aim is to test whether any collective effects may appear in the case of a larger number of components, which are not present in binary mixtures. In order to make these studies well-defined, reasonably comprehensive, and also going beyond extensively studied balanced mixtures, we consider the smallest possible extension of a two-component fermionic mixture by introducing additional interaction with a single third-component particle. We aim to investigate if and how interactions with this impurity change the internal properties of the mixture. Moreover, we inspect non-classical correlations this may induce. In this way, we also link our work to the well-studied problem of a single impurity interacting with several spin-polarized fermions Wenz et al. (2013); Włodzyński (2022).
Our work is organized as follows. In Sec. II we introduce the theoretical model of the system studied and we describe its Hamiltonian in the convenient formalism. Then in Sec. III we briefly explain the exact diagonalization method in the energy cut-off framework. In Sec. IV we analyze single- and two-particle density distributions and we identify a specific transition in the many-body ground state of the system induced by interactions with impurity. In Sec. V we study this transition in detail and we show that it has intriguing universal properties that do not depend essentially on either the number of particles in the system or the internal interactions in the mixture. Finally, we conclude in Sec. VI.
@@ II The system
In our work, we focus on the simplest possible extension of the two-component mixture of fermions to scenarios where interactions with an additional third-component particle (impurity) are present. We assume that all the particles have the same mass $m$, they are confined in the same one-dimensional parabolic trap of frequency $\Omega$, and they mutually interact via zero-range contact interactions in the $s$-wave channel. This implies, of course, that, due to the Pauli exclusion principle, only interactions between particles from different components do not vanish. Since in our considerations the third component is essentially distinguished, we write the Hamiltonian of the system conveniently as
$\hat{\cal H}=\hat{\cal H}_{AB}+\hat{\cal H}_{C}+\hat{\cal H}_{I}.$ (1)
In this sum, the Hamiltonian $\hat{\cal H}_{AB}$ describes the two-component mixture of fermions and in the second quantization reads
$\hat{\cal H}_{AB}$ $=\int\!\mathrm{d}x\sum_{\sigma}\,\hat{\Psi}_{\sigma}^{\dagger}(x)\left[-\frac{\hbar^{2}}{2m}\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}+\frac{m\Omega^{2}}{2}x^{2}\right]\hat{\Psi}_{\sigma}(x)$
@@
Here $\sigma\in\{A,B\}$ enumerates mixture’s components while $\hat{\Psi}_{\sigma}(x)$ is a field operator annihilating $\sigma$-component particle at position $x$. These operators obey fermionic anti-comutation relations $\{\Psi_{\sigma}(x),\Psi^{\dagger}_{\sigma^{\prime}}(x^{\prime})\}=\delta_{\sigma\sigma^{\prime}}\delta(x-x^{\prime})$ and $\{\Psi_{\sigma}(x),\Psi_{\sigma^{\prime}}(x^{\prime})\}=0$. The dimensionless parameter $g_{0}$ quantifies an effective interaction strength between components.
Contrary, the Hamiltonian $\hat{\cal H}_{C}$ describes a third-component particle confined in the same trap. In the first quantization, it is represented by the differential operator of the form
$\hat{\cal H}_{C}=-\frac{\hbar^{2}}{2m}\frac{\mathrm{d}^{2}}{\mathrm{d}z^{2}}+\frac{m\Omega^{2}}{2}z^{2},$ (2b)
@@
Finally, the Hamiltonian $\hat{\cal H}_{I}$ is responsible for interactions between the two-component mixture and the third particle. It can be written as
$\hat{\cal H}_{I}=g\sqrt{\frac{\hbar^{3}\Omega}{m}}\int\!\mathrm{d}x\sum_{\sigma}\,\hat{\Psi}^{\dagger}_{\sigma}(x)\delta(x-z)\hat{\Psi}_{\sigma}(x),$ (2c)
@@ III The Method
where $|\phi_{n}\rangle$ is a state of the third-component particle occupying $n$-th single-particle orbital, while two integer vectors $\vec{\alpha}=(\alpha_{1},\ldots,\alpha_{N})$ and $\vec{\beta}=(\beta_{1},\ldots,\beta_{N})$ encode single-particle orbitals occupied by individual fermions from components $A$ and $B$, respectively. Indistinguishability and fermionic nature of particles is guaranteed by imposing additional constrains of the form $\alpha_{1}>\alpha_{2}>\ldots>\alpha_{N}$ and $\beta_{1}>\beta_{2}>\ldots>\beta_{N}$. States (5) are obviously eigenstates of the non-interacting Hamiltonian ($g_{0}=g=0$) and have energies $E_{\vec{\alpha},\vec{\beta},n}=\hbar\Omega\left[(2N+1)/2+n+\sum_{i}(\alpha_{i}+\beta_{i})\right]$. Non-interacting ground state has energy $E_{0}=\hbar\Omega(N^{2}+1/2)$.
@@
To diagonalize numerically the Hamiltonian one needs to cut the basis. In our approach, we do not perform simple cuts on single-particle excitations but rather we limit the many-body basis to states having total (non-interacting) energy $E_{\vec{\alpha}\vec{\beta}n}$ not larger than some fixed energy limit $E_{C}$. In this way, we do not exclude from the description low-energy states having only a few highly excited particles, while we neglect high-energy states with many low-excited particles. It was shown in previous studies that such a method gives a much better approximation of interacting many-body states with the same usage of numerical resources Haugset and Haugerud (1998); Płodzień et al. (2018); Chrostowski and Sowiński (2019); Rojo-Francàs et al. (2022). In our case, determination of the appropriate value of cut-off energies $E_{C}$ for specific systems considered is based on a convergence test of the many-body ground state. A detailed description of this procedure and the cut-off energy values used is provided in the Appendix A. Following this procedure, for a given number of particles $N$ and interaction strengths $g_{0}$ and $g$, we represent the Hamiltonian (1) in the truncated many-body basis and diagonalize with the implicitly-restarted Arnoldi-Lanczos algorithm Lehoucq et al. (1998). As a result, we obtain the lowest many-body eigenstates represented as a superposition of basis states $|{\vec{\alpha},\vec{\beta},n}\rangle$ and their energies. In this way, we have particular access to a whole structure of the many-body ground state $|\mathtt{G}\rangle$. We present exemplary spectra for the system of five particles ($N=2$) as a function of the interaction strengths in Fig. 1. Let us mention here that in the case of harmonic confinement, further optimization techniques of the many-body basis are possible Volosniev, A.G. et al. (2015); Bellotti, Filipe F. et al. (2017); Kościk (2018, 2020). They are vital when large interaction strengths are considered. In the cases of relatively weak interactions studied in our work, this kind of optimization has limited importance. Therefore, we do not introduce them to our scheme.
Figure 1: Low-energy spectrum of the many-body Hamiltonian (1) obtained for a system containing an impurity interacting with the two-component mixture of four particles ($N=2$). (a) Eigenenergies as functions of interactions with the impurity $g$ for a fixed internal interaction $g_{0}$. (b) Eigenenergies as functions of internal interactions $g_{0}$ for a fixed interactions with the impurity $g$. In all plots, energies are expressed in natural units of energy $\hbar\Omega$, and for clarity, we shift them by the ground-state energy at $g=0$ for (a) and $g_{0}=0$ for (b), respectively.
@@ IV Impact of the impurity
Figure 2: Impact of interactions with the third-component particle on the two-component mixture containing four particles ($N=2$). Different columns correspond to different values of interaction strength with the impurity $g$. In all plots, inter-component interaction strength is fixed, $g_{0}=1$. (a) Single-particle density profiles of the mixture’s components $n_{\sigma}(x)$ (thick blue) and the third-component particle $m(x)$ (thin red). For a better comparison, the non-interacting distribution of the mixture’s components is additionally displayed with a light blue shadow. (b) Single-particle reduced density matrix of the impurity $\rho(x,x^{\prime})$. (c) Two-particle density profile $\mu_{\sigma}(x,z)$. (d) Distribution of correlations ${\cal G}_{\sigma}(x,z)$ between impurity and particle from the component $\sigma$. Note, that in the left plot, we consider the interaction strength $g$ slightly positive ($g=0.01$) since for vanishing interactions the distribution vanishes. In this case, the distribution intensity is also scaled by a factor $10^{2}$ to make it visible. The upper and lower values on the color bar refer to plots (c) and (d), respectively. In all plots, positions are given in units of $a_{0}$. In panels (a) and (b), densities are expressed in units of $a_{0}^{-1}$, while in panels (c)-(d), in units of $a_{0}^{-2}$.
Figure 3: Distribution of internal two-particle correlations ${\cal K}(x,y)$ in the system containing impurity interacting with the two-component mixture containing four particles ($N=2$) for different internal interaction strengths $g_{0}$ and interactions with the impurity $g$. For any fixed interaction $g_{0}$, along with increasing repulsion $g$, the transition between two different regimes is visible. In all plots, positions are expressed in units of $a_{0}$ while densities in units of $a_{0}^{-2}$.
@@
where $\hat{\mathbb{P}}_{x}=|x\rangle\langle x|$ projects to an eigenstate $|x\rangle$ of position operator. Of course, due to the exchange symmetry $A\leftrightarrow B$ explained before, both density distributions $n_{A}(x)$ and $n_{B}(x)$ are exactly the same. In Fig. 2(a), we present these distributions (thick blue and thin red lines) for the system of five particles ($N=2$) for different interaction strengths $g$, keeping fixed internal interaction $g_{0}=1$. For better comparison, we additionally display with the shadowed area the non-interacting ($g_{0}=g=0$) distribution $n_{\sigma}(x)$. It is clear that along with increasing $g$, the distribution of the impurity $m(x)$ significantly changes its character and for sufficiently large repulsions it is split and pushed out from the center where the mixture is concentrated. In contrast, distributions of mixture components $n_{\sigma}(x)$ are only slightly affected. Since the impact of interactions on the impurity is significant, it is also worth analyzing the off-diagonal order of the separated particle. This is encoded directly in the single-particle reduced density matrix which in our notation can be written as
$\rho(x,x^{\prime})=\langle\mathtt{G}|\left(|x\rangle\langle x^{\prime}|\otimes\mathbb{1}\right)|\mathtt{G}\rangle.$ (7)
Obviously, a diagonal part of this quantity is equal to a single-particle density, $m(x)=\rho(x,x)$. As clearly seen in Fig. 2(b), the reduced density matrix undergoes a substantial change when the system is driven between the two regimes.
This behavior of single-particle properties can be investigated further by considering the two-particle density distribution
@@
which encodes the joint probability of finding the impurity at position $z$ and one of the particles from the component $\sigma$ at position $x$. The distribution for the same parameters of the system is shown in Fig. 2(c). In addition to the spatial separation already deduced from single-particle distributions, the two-particle density clearly indicates that some correlations between mixture and impurity are developed. Most significantly, they are visible along the diagonal direction $z=x$, i.e., the probability of finding both particles at the same position quickly vanishes with increasing interaction $g$. Of course, this effect cannot be captured by single-particle distributions and is not present in the direct product $n_{A}(x)m(z)$. To make it evident, we additionally display in Fig. 2(d) the distribution of correlations defined as
@@
From the perspective of the internal properties of the mixture, the description is slightly simpler. As noticed already, interactions with the impurity do not significantly influence the single-particle distribution of the mixture. It turns out that this is also the case when one considers the inter-component two-particle distribution
@@
The impact of the impurity is slightly noticeable when the distribution of two-particle correlations is analyzed. It is defined analogously to correlations (9) as
@@
In Fig. 3 we show this distribution for different values of interactions $g_{0}$ and $g$ for the same system of five particles ($N=2$). The magnitude of correlations is very small when compared to the magnitude of correlations in ${\cal G}_{\sigma}(x,z)$ and the transition is not as clearly visible as in Fig. 2. However, one can still notice that correlations along the anti-diagonal (especially far from the center) change their character when the system undergoes the transition. Nonetheless, it is true that from the perspective of internal correlations within the mixture, the structural transition is not very clearly manifested.
@@
Figure 4: Impact of interactions with the third-component particle on the two-component mixture containing four particles ($N=2$) when the system is confined in a rectangular potential well (12). Different columns correspond to different values of interaction strength with the impurity $g$. In all plots, inter-component interaction strength is fixed, $g_{0}=1$. (a) Single-particle density profiles of the mixture’s components $n_{\sigma}(x)$ (thick blue) and the third-component particle $m(x)$ (thin red). (b) Single-particle reduced density matrix of the impurity $\rho(x,x^{\prime})$. Note that for sufficiently large interaction $g$, the impurity is separated similarly as in the case of a harmonic confinement. In all plots, positions are expressed in the width of the well $L$, while densities are in $L^{-1}$.
@@
From phenomenological argumentation, some kind of separation of the components is suspected. When interactions are strong enough, the single-particle excitations become energetically favourable over the interaction energy cost. Then, to reduce the overlap of the densities, the system excites particles. It should be noted, however, that the separation observed here is substantially different from separations observed in systems of single-component fermions interacting with the impurity Lindgren et al. (2014). It was shown that in those systems, the impurity remains in the middle of the trap while the fermionic component is (only slightly) split and pushed out from the center. In our case, although repulsive interactions in the fermionic mixture are present, the distribution of the impurity is split and driven to the edges. As noted in the literature, in the case of the single-component Fermi sea, the existence of this kind of separation requires a noticeable mass difference between impurity and other particles Pȩcak et al. (2016); Włodzyński (2022). Still, however, there is a substantial difference between the scenario studied here and two-component mass-imbalanced systems. To demonstrate this, it is sufficient to examine how the separation depends on the shape of the external potential, when the Hamiltonian (1) has the following form
$\hat{\cal H}_{AB}$ $=\int_{-L}^{L}\!\mathrm{d}x\sum_{\sigma}\,\hat{\Psi}_{\sigma}^{\dagger}(x)\left(-\frac{\hbar^{2}}{2m}\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}\right)\hat{\Psi}_{\sigma}(x)$
$+g_{0}\frac{\hbar^{2}}{mL}\int_{-L}^{L}\!\mathrm{d}x\,\hat{\Psi}^{\dagger}_{B}(x)\hat{\Psi}^{\dagger}_{A}(x)\hat{\Psi}_{A}(x)\hat{\Psi}_{B}(x),$ (12a)
$\hat{\cal H}_{C}$ $=-\frac{\hbar^{2}}{2m}\frac{\mathrm{d}^{2}}{\mathrm{d}z^{2}}+V(z),$ (12b)
$\hat{\cal H}_{I}$ $=g\frac{\hbar^{2}}{mL}\int_{-L}^{L}\!\mathrm{d}x\sum_{\sigma}\,\hat{\Psi}^{\dagger}_{\sigma}(x)\delta(x-z)\hat{\Psi}_{\sigma}(x),$ (12c)
where $V(z)=0$ for $|z|<L$ and $V(z)\rightarrow\infty$ otherwise. It is known that in the case of two-component systems with different masses, the lighter component undergoes separation in the harmonic potential, but in a rectangular potential well, the heavier component is separated Pęcak and Sowiński (2016). In Fig. 4, we show that in the case of a two-component mixture interacting with the impurity described with Hamiltonian (12), there is no such change induced by the shape of the potential – the impurity density is again subjected to separation.
@@
Figure 5: Properties of the fidelity susceptibility for different numbers of particles in the system. (a) The fidelity susceptibility $\chi(g_{0};g)$ as a function of interaction $g$ for different internal interactions $g_{0}$ (legend in panel (c)). (b) Position of the transition $G_{N}(g_{0})$ as a function of internal interaction $g_{0}$. (c) Susceptibilities from the left plot after rescaling in magnitude by $\chi_{0}(g_{0})$ and centering around the transition interaction $G_{N}(g_{0})$. All plots collapse to the same universal curve, signalling the specific universality of the transition. (d) Numerical collapse to the universal curve $\Phi$ (dashed black) of all the fidelity susceptibility curves (different points and colors) after applying appropriate scaling (16). Note that deviations from the universality are diminished when the number of particles in the mixture increases.